A uniform pressure P is exerted on all sides of a solid cube at a temp t0C.Given the coeff. of volume expansion of cube is  and bulk modulus of elasticity B, show that the amount of temperature of the cube to be raised in order to bring back its original volume (before pressure was applied)  = P/B.


Grade:9

1 Answers

Shubham Patil
6 years ago
Bulk modulus=kCoefficient of expansion=aB=dP/(-dV/V)k=-VdP/dV-dV/V=(0-P)/kdV/V=P/kNow dV is change in volume i.e. a∆ta∆t=P/k∆t=P/KaHope it helps!!!